OpenStudy (anonymous):

W varies jointly with the inverse of the cube root h and directly with A; W = 5 when h = 216 and A = 200

OpenStudy (anonymous):

Could someone please help me?

OpenStudy (anonymous):

btw is there any key words i should know

OpenStudy (anonymous):

\( W=\dfrac{A}{\sqrt[3]{h}}\)

OpenStudy (anonymous):

I think they want you symbolize the word problem

OpenStudy (anonymous):

so W = 3A / 20 3sqrt(h)?

OpenStudy (anonymous):

your answer isn't on here

OpenStudy (anonymous):

never mind

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

would you mind helping with one more?

OpenStudy (mathmate):

@avonniculleyy The question does not seem to be complete (didn't tell you what it expects).

OpenStudy (anonymous):

M varies jointly with the inverse of the square of B and directly with the cube of b.

OpenStudy (anonymous):

its telling me to describe the variation

OpenStudy (anonymous):

@OOOPS

OpenStudy (mathmate):

If they want the equation for W, there needs to be a proportional constant, something like C or K, because it says "M varies jointly... ", didn't say equal.

OpenStudy (mathmate):

A varies directly with B is written A=kB

OpenStudy (mathmate):

Is this your question, or you have another one?

OpenStudy (anonymous):

I think its, M = kB^2/b^3

OpenStudy (anonymous):

that was my question

OpenStudy (mathmate):

What about it? M=kB^2/b^3 is not related to the earlier question, is it?

OpenStudy (anonymous):

yea it is

OpenStudy (anonymous):

its one of the answer

OpenStudy (mathmate):

I don't see B and b in the previous question!

OpenStudy (anonymous):

M varies jointly with the inverse of the square of B and directly with the cube of b.

OpenStudy (mathmate):

Ah, that's different! :)

OpenStudy (anonymous):

This is my last question and I want to get it right

OpenStudy (mathmate):

``` M varies jointly with the inverse of the square of B and directly with the cube of b. ``` M varies, so M=k(...), ok so far?

OpenStudy (mathmate):

X varies directly with the square of Y is written X=kY^2.

OpenStudy (anonymous):

So it's M = kB^2 / b^3?

OpenStudy (mathmate):

You have to check your answer against the question for "Directly" and "inversely" .

OpenStudy (mathmate):

X varies inversely with the cube of Z is written X=k/Z^3.

OpenStudy (mathmate):

So if we have X varies directly with the square of Y and inversely with the cube of Z is written X=kY^2/Z^3. (remember, this is an example, not your question).

OpenStudy (anonymous):

ohh. okay

OpenStudy (mathmate):

So what is your final answer?

OpenStudy (anonymous):

do you mind staying here 4 a bit? and check my answer?

OpenStudy (mathmate):

no problem. Work it out and post!

OpenStudy (anonymous):

Will do!

OpenStudy (anonymous):

is there for me 2 add u so I can text u when I need your help?

OpenStudy (anonymous):

are you there @mathmate

OpenStudy (mathmate):

You can message me on OS, or tag me when I'm online. Unfortunately I am on and off a lot, so message is probably more efficient.

OpenStudy (anonymous):

Write a general formula to describe the variation: x varies jointly with the inverse of r and the inverse of the square of the sum of y and z.

OpenStudy (mathmate):

what did you get for the previous one?

OpenStudy (anonymous):

I got M = k / (B+b)^2

OpenStudy (mathmate):

The question didn't say the "sum of B and b", so it cannot be right. Think of the following scheme: